Two lamps rated 60w and 240v are connected in series. What is the total power dissipated in both.
1
Expert's answer
2019-07-29T08:25:16-0400

The resistance of one lamp

"R_1=\frac{V^2}{P_1}=\frac{(240\:\rm{V})^2}{60\:\rm{W}}=960\:\Omega"

The resistance of two lamps connected in series

"R=2R_1=2\times 960\:\Omega=1920\:\Omega"

Finally, the total power dissipated in both lamps is

"P=\frac{V^2}{R}=\frac{(240\:\rm{V})^2}{1920\:\Omega}=30\:\rm{W}"